determine the best level of output if a firm operates in a perfectly competitive industry where p=#8 and TC=Q3-5Q2+10Q+4
1
Expert's answer
2020-09-15T10:10:32-0400

The best level of output is at MC = P.

"MC = TC'(Q) = 3Q^2 - 10Q + 10,"

"3Q^2 - 10Q + 10 = 8,"

"3Q^2 - 10Q + 2 = 0,"

"Q1 = (10 + 8.72)\/6 = 3.12,"

"Q2 = (10 - 8.72)\/6 = 0.21."